# Functoriality of Odd and Generalized Khovanov Homology in $\mathbb{R}^{3}\times I$

### v1

@@ 2.5   Type X and Type Y Configurations
From the convention in 3, there emerges a pair of exceptional arrangements which, a priori, are neither commuting, anticommuting, nor $\pi$-commuting as they are annihilated by 2.
Figure 2.2. Cobordisms of type X and type Y configurations
In order to define an generalized Khovanov theory, one must artificially decide that one of these configurations commutes and that the other $\pi$-commutes. These two different generalized Khovanov theories are called type X or type Y based on which pair of cobordisms in Figure 2.2 $\pi$-commutes. It was first observed in [ORS2007] that either choice produces the same final invariant, although their proof was incorrect. Putyra attempted in [Pu2015] to provide a corrected proof, but his proof also appears to be incorrect. This is Lemma 12, which we will prove in Section 3. We will use type Y sign assignments following Putyra in [Pu2015].
@@ 3   Generalized Khovanov Theory
Oriented links allow us to easily assign a direction to crossings (although it can still be defined for unoriented links using a variation on the right-hand rule). This convention is shown in Figure 3.1
Figure 3.1. Positive and negative crossings for directed links
We will now construct Putyra’s version of odd Khovanov homology. Then we will review portions of the proof of invariance, as they will be needed to extend odd Khovanov homology to link cobordisms and to prove that it is functorial. We will end this section by defining the chain maps that are assigned to each of the elementary cobordisms in $\mathcal{C}ob^{4}$.
3.1   Links with Oriented Crossings
Odd Khovanov homology is not computed from a simple link diagram, but from a link diagram that has been enhanced with orientations on the crossings. Each crossing has two possible orientations, which are displayed in Figure 3.2.
Figure 3.2. The two possible orientations on a crossing
Note that the arrows specifying the crossing orientation are chosen so that they connect the regions that lie to the left if one approaches the crossing along the overstrand from either side.
@@ 3.2   Crossing Resolution
A link diagram with $n$ crossings gives rise to $2^{n}$ planar diagrams corresponding to all possible combinations of replacing each crossing with the vertical or horizontal resolution, where the terms “vertical” and “horizontal” refer to the pictures in Figure 3.2. The rule for resolving a crossing is shown in Figure 3.3. If the crossings are labeled from 1 to $n$, then each diagram, $D_{\alpha}$, can be assigned a binary label $\alpha\in\{0,1\}^{n}$ with a 0 in the $i$th place if the $i$th crossing was replaced with the vertical resolution, and a 1 if it was replaced with the horizontal resolution.
Figure 3.3. The vertical and horizontal resolution of a crossing
@@
To denote that an index is in superposition we use a $\star$ at that index.
We assign a cobordism $d_{{\dots}\star{\dots}}:D_{{\dots}0{\dots}}\rightarrow D_{{\dots}1{\dots}}$ to each pair of planar diagrams that differ by a single resolution. The cobordism $d_{{\dots}\star{\dots}}$ is a morphism in $\mathcal{C}ob^{3}$ whose initial frame is the vertical resolution and whose terminal frame is the horizontal resolution with a saddle between (as depicted in Figure 3.3). The saddle inherits its orientation from the resolved crossing’s orientation.
@@
The degree of a resolution is the number of horizontal resolutions in the resulting diagram.
@@ 3.2.1   The Generalized Khovanov Bracket
The cube of resolutions of a link diagram $D$ with $n$ crossings is an $n$ dimensional cube with the $2^{n}$ planar diagrams $D_{\alpha}$ for $\alpha\in\{0,1\}^{n}$ as its vertices and the corresponding cobordisms as its edges.
Each 2-dimensional face of the cube of resolutions corresponds to two vertices living in the superposition of their resolution, while all others are resolved. In order to define generalized Khovanov homology we need a chain complex. Therefore, all faces must anticommute so that the differential squares to zero. Each face falls into one of a few basic types and has an inherited sign $\sigma_{i,j}$ of either 1 or $\pi$ corresponding with the indices of the non-resolved crossings in $D_{\dots\star\dots\star\dots}$. The inherited signs appear following diagrams considered up to planar isotopy. Each tangle diagram can be closed with any crossingless tangle that connects oppositely decorated endpoints.
Figure 3.4. Commutivity of faces for type $Y$ sign assignments
  Commuting Faces (Group C) — $\sigma_{i,j}=1$
@@
Refer to caption Refer to caption .   $\pi$-commuting Faces (Group $\pi$) — $\sigma_{i,j}=\pi$
@@
Refer to caption or Refer to caption   X Face (Group X) — $\sigma_{i,j}=1$
Refer to caption or Refer to caption   Y Face (Group Y) — $\sigma_{i,j}=\pi$
@@
We need to make an assignment of a sign $\epsilon_{\dots\star\dots}$ to each edge such that
(31) $\prod_{\{\iota_{i},\iota_{j}\}\in\{\{0,{\star}\},\{1,{\star}\},\{{\star},0\},\{{\star},1\}\}}\epsilon_{\dots\iota_{i}\dots\iota_{j}\dots}=-\sigma_{i,j}$
Such a sign assignment can always be made ([ORS2007],[Pu2015]), but it is not unique.
@@
The generalized Khovanov cube is the cube of resolutions with each edge multiplied by its assigned sign. [^6] The odd/even Khovanov cubes normally refer to the cubes of resolutions built over the relevant algebra rather than the cube of resolution in the Bar-Natan category.
Definition 8.
The generalized Khovanov bracket is the grading-shifted flattening of the generalized Khovanov cube created by taking the direct sum of all vertices with the same degree.
The generalized Khovanov bracket is a chain complex, as all maps in the Khovanov cube raise the degree by one, and before collapsing the cube we ensured that all faces anticommute. The grading shift in the generalized Khovanov bracket is determined by the number of negative crossings in the link diagram. Note that we interpret the generalized Khovanov bracket as an object of the category $\textnormal{Kom}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen}))$.
@@
The unsigned cobordisms in the generalized Khovanov cube will be denoted with $d$ while the maps in the generalized Khovanov bracket are denoted with $\partial$.
@@
For a link diagram $L$ the generalized Khovanov bracket is denoted $\llbracket L\rrbracket$, additionally, the specific $\alpha$th vertex is denoted by $\llbracket L\rrbracket_{\alpha}$ so that $\llbracket L\rrbracket_{\alpha}=D_{\alpha}$.
@@ 3.3   Odd and Generalized Khovanov TQFTs
We can construct odd or even Khovanov homology by starting with the generalized Khovanov bracket, setting $\pi$ to $-1$ or $1$, and applying the appropriate chronological TQFT functor [Pu2015]. Explicitly, the odd chronological TQFT functor can be defined as in [ORS2007] and sends a resolution $R$ to the exterior algebra
@@
where $V(R)$ denotes the free $\mathbb{Z}$-module formally generated by the connected components of $R$. The map assigned to a planar saddle cobordism that merges two components $R_{1}$ and $R_{2}$ of $R$ is defined via the obvious quotient map $F(R)\rightarrow F(R)/(R_{1}=R_{2})$, whereas the map assigned to a split saddle is defined via the map $F(R)/(R_{1}=R_{2})\rightarrow F(R)$ induced by left-multiplication by $R_{1}-R_{2}$ (see [ORS2007] for details). There is a generalized Khovanov TQFT functor that is usually defined in the language of [Bn2002], but that can be defined in analogy $F(R)$ by replacing the exterior algebra by the noncommutative polynomial ring over $𝕜\Bbbk$ in the components of $R$, modulo the relations
@@
where $R_{1}$ and $R_{2}$ are components of $R$. These relations also imply
@@
and so the above map $F(R)/(R_{1}=R_{2})\rightarrow F(R)$ induced by left-multiplication by $R_{1}-R_{2}$ extends to the generalized setting if one instead multiplies by $R_{1}+\pi R_{2}$. In fact, the resulting map for a planar split saddle cobordism corresponds precisely to the generalized split map from [Pu2015].

### v2

@@ 2.5   Type X and Type Y Configurations
From the diamond relation, there emerges a pair of exceptional arrangements which are both commuting and $\pi$-commuting because the involved cobordisms are annihilated by $1-\pi$. These two configurations are called type X and type Y and are shown in Figure 3.
Figure 3. Cobordisms of type X and type Y configurations
Note that Figure 3 correctly reflects the embeddings of the shown cobordisms into the oriented manifold $S^{2}\times I\supset\mathbb{R}^{2}\times I$. In fact, the two types can be transformed into each other by applying an orientation-reversing homeomorphism of the plane $\mathbb{R}^{2}$.
In order to define a generalized Khovanov theory, one must artificially decide that one of these configurations commutes and that the other $\pi$-commutes. By consequence, there are two different generalized Khovanov theories, which are called type X or type Y based on which pair of cobordisms in Figure 3 $\pi$-commutes. It was first observed in [ORS2007] that either choice produces the same final invariant, although the proof given there was incorrect. Putyra in [Pu2015] outlined a corrected proof, but his argument is not clear to us and does not address the functoriality of the type X and type Y theories under link cobordisms. This will be done in Lemma 13 in Subection 3.8, where we will show that the type X and type Y theories provide naturally isomorphic functors. For the most part of this paper, we will use type Y sign assignments following Putyra [Pu2015].
@@ 3   Generalized Khovanov Theory
We will now construct the generalized Khovanov bracket of an oriented link diagram $D$. Then we will review portions of the proof of invariance, as they will be needed to extend the generalized Khovanov bracket to link cobordisms and to prove that it is functorial. We will end this section by defining the chain maps that are assigned to each of the elementary cobordisms in $\mathcal{C}ob^{4}$.
The generalized Khovanov bracket is not constructed from a simple oriented link diagram, but from an oriented link diagram that has been enhanced with a local orientation at each of its crossings. Each crossing has two possible orientations, which are displayed in Figure 4 and represented by arrows.
Figure 4. The two possible orientations on a crossing
Note that the arrows specifying the crossing orientation are chosen so that they connect the regions that lie to the left if one approaches the crossing along the overstrand from either side. In the following, we will assume that link diagrams are oriented and equipped with orientations on their crossings, unless otherwise stated.
@@ 3.1   Crossing Resolution
A link diagram $D$ with $n$ crossings gives rise to $2^{n}$ planar diagrams corresponding to all possible combinations of replacing each crossing with the vertical 00-resolution or the horizontal $1$-resolution. Here the terms “vertical” and “horizontal” refer to the pictures in Figures 4 and 5. In intrinsic terms, the 00-resolution is the resolution whose arcs veer to the left as one approaches the original crossing along the overstrand from either side, and the $1$-resolution is the resolution whose arcs veer to the right.
If the crossings of $D$ are labeled from 1 to $n$, then each resolved diagram, $D_{\alpha}$, can be assigned a binary label $\alpha\in\{0,1\}^{n}$ with a 0 in the $i$th place if the $i$th crossing was replaced with the 00-resolution, and a 1 if it was replaced with the $1$-resolution.
Figure 5. The two resolutions of a crossing
@@
To denote that a digit in a binary string is in a “superposition” of a 00 and a $1$, we use a $\star$ for that digit.
We assign a cobordism $d_{{\dots}\star{\dots}}:D_{{\dots}0{\dots}}\rightarrow D_{{\dots}1{\dots}}$ to each pair of resolved diagrams that differ at a single crossing $c$ of $D$. Specifically, $d_{{\dots}\star{\dots}}$ is a saddle cobordism from the diagram $D_{{\dots}0{\dots}}$, in which $c$ is replaced by a 00-resolution, the diagram $D_{{\dots}1{\dots}}$, in which $c$ is replaced by a 1-resolution (as depicted in Figure 5). The saddle point in $d_{{\dots}\star{\dots}}$ inherits a local orientation from the orientation of the resolved crossing $c$, and thus $d_{{\dots}\star{\dots}}$ be viewed as a morphism in $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$.
@@
The degree of a resolution is the number of $1$-resolutions in the resulting diagram.
@@ 3.1.1   The Generalized Khovanov Bracket
The cube of resolutions of a link diagram $D$ with $n$ crossings is an $n$-dimensional cube with the $2^{n}$ planar diagrams $D_{\alpha}$ for $\alpha\in\{0,1\}^{n}$ as its vertices and the corresponding cobordisms as its edges.
Each 2-dimensional face of the cube of resolutions corresponds to a sequence in $\{0,1,\star\}^{n}$ with exactly two $\star$’s, and thus corresponds to an (unoriented) link diagram $D_{\dots\star\dots\star\dots}$ in which all but two of the crossings of $D$ are resolved. In order to define a chain complex, we must make sure that all 2-dimensional faces of the cube of resolutions anticommute, so that the resulting differential squares to zero.
Each 2-dimensional face falls in one of a few basic types, which are shown in Figure 6. In this figure, we are actually depicting the 2-crossing link diagrams $D_{\dots\star\dots\star\dots}$ corresponding to 2-dimensional faces. In the cases where we have drawn a tangle diagram instead of a link diagram, the convention is that the tangle diagram can be closed up with any crossingless tangle that connects oppositely decorated endpoints. In diagrams where we have drawn crossing orientations, it is understood that one either has to choose the single arrows at both crossings, or the double arrows at both crossings. Moreover, all link diagrams in Figure 6 are to be considered up to planar isotopy. As indicated in the figure, each face has an element $\sigma_{\dots\star\dots\star\dots}\in\{1,\pi\}$ assigned to it that depends on whether the face is naively commuting or $\pi$-commuting.
Figure 6. Commutativity of faces for type $Y$ sign assignments
  Commuting Faces (Group C) — $\sigma_{\dots\star\dots\star\dots}=1$
@@
Refer to caption Refer to caption .   $\pi$-commuting Faces (Group $\pi$) — $\sigma_{\dots\star\dots\star\dots}=\pi$
@@
Refer to caption or Refer to caption   X Face (Group X) — $\sigma_{\dots\star\dots\star\dots}=1$
Refer to caption or Refer to caption   Y Face (Group Y) — $\sigma_{\dots\star\dots\star\dots}=\pi$
@@
Note that in Figure 6, we have depicted the conventions for type Y assignments, where faces of type X are considered to be naively commuting, while faces of type Y are considered to be naively $\pi$-commuting. For type X assignments, the latter conventions would be reversed.
[…]
(31) $\prod_{\{\alpha_{i},\alpha_{j}\}\in\{\{0,{\star}\},\{1,{\star}\},\{{\star},0\},\{{\star},1\}\}}\epsilon_{\dots\alpha_{i}\dots\alpha_{j}\dots}=-\sigma_{\dots\star\dots\star\dots}$
Such a sign assignment can always be made ([ORS2007],[Pu2015]), but it is not unique. This will be discussed in more detail in Subsection 3.5 below.
@@
The generalized Khovanov cube of $D$ is the cube of resolutions with each edge map $d_{\dots\star\dots}$ multiplied by its assigned element $\epsilon_{\dots\star\dots}\in\{\pm 1,\pm\pi\}$.
To obtain the generalized Khovanov bracket, we now “flatten” the generalized Khovanov cube to a chain complex. The $i$th chain object in this complex is the formal direct sum
$\llbracket D\rrbracket^{i}:=\bigoplus_{\deg(\alpha)=i+n_{-}}D_{\alpha},$
viewed as an object of $\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen})$, where $n_{-}$ denotes the number of negative crossings in $D$ (defined as in Figure 7). The $i$th differential $\partial^{i}\colon\llbracket D\rrbracket^{i}\rightarrow\llbracket D\rrbracket^{i+1}$ is given by all edge maps $\epsilon_{\dots\star\dots}d_{\dots\star\dots}\colon D_{\alpha}\rightarrow D_{\beta}$ in the generalized Khovanov cube that start at a vertex $\alpha$ with $\deg(\alpha)=i+n_{-}$. Note that we interpret the resulting generalized Khovanov bracket as an object of the category $\textnormal{Kom}(\textnormal{Mat}(\mathcal{C}hron\mathcal{C}ob^{3}_{gen}))$.
Figure 7. Positive and negative crossings
@@
The underlying saddle cobordisms in the generalized Khovanov cube will be denoted with $d$ while the edge maps $\epsilon_{\ldots{\star}\ldots}d_{\ldots{\star}\ldots}$ will be denoted with $\partial$.
@@
For an oriented link diagram $D$ with chosen crossing orientations, the generalized Khovanov bracket will be denoted $\llbracket D\rrbracket$. Additionally, the diagram at vertex $\alpha$ of the cube of resolutions will be denoted by $\llbracket D\rrbracket_{\alpha}$, so that $\llbracket D\rrbracket_{\alpha}:=D_{\alpha}$.
Note that since all of the $d_{\ldots{\star}\ldots}$ are saddle cobordisms, the differential $\partial^{i}$ in the generalized Khovanov bracket a priori has quantum degree $-1$. One can achieve that each $\partial^{i}$ has quantum degree zero by formally raising the quantum grading of each $\llbracket D\rrbracket^{i}$ by $i+n_{+}-n_{-}=i+w(D)$ where $n_{-}$ denotes the number of negative crossings (as before), $n_{+}$ denotes the number of positive crossings, and $w(D):=n_{+}-n_{-}$ denotes the writhe of $D$. This construction is described in detail in [Bn2005] and also ensures that the chain map induced by a link cobordism is homogeneous of quantum degree given by the Euler characteristic of the link cobordism. In this paper, we will not pursue this construction any further, but we will describe an analogous construction, for the supergrading, in the next subsection.
3.2   Supergraded Refinement
Recall that a $𝕜\Bbbk$-linear category is called a supercategory if its morphism sets are $\mathbb{Z}_{2}$-graded $𝕜\Bbbk$-modules such that the $\mathbb{Z}_{2}$-grading is additive under composition of morphisms. A supercategory $\mathcal{C}$ can be extended to a “refined” supercategory $\mathcal{C}^{r}$ in which objects come with formal grading shifts. The objects in the refined supercategory $\mathcal{C}^{r}$ are symbols $x\{m\}$ with $x\in\mathcal{C}$ and $m\in\mathbb{Z}_{2}$, and the morphism sets are given by
$\operatorname{Hom}(x\{m\},y\{n\}):=\operatorname{Hom}(x,y),$
but with the $\mathbb{Z}_{2}$-grading shifted by $n-m$. Thus, if a morphism $f$ has degree $d\in\mathbb{Z}_{2}$ in $\operatorname{Hom}(x,y)$, then the same morphism has degree $d+n-m$ when viewed as a morphism in $\operatorname{Hom}(x\{m\},y\{n\})$.
The category $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$ can be viewed as a supercategory with $\mathbb{Z}_{2}$-grading given by the superdegrading from Subsection 2.2. Working in the refinement of $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$, we can now define a refined generalized Khovanov bracket. This refined Khovanov bracket is obtained by replacing each $\llbracket D\rrbracket_{\alpha}=D_{\alpha}$ by $D_{\alpha}\{s_{\alpha}\}$, where $s_{\alpha}$ is the modulo 2 reduction of
(32) $S(D,\alpha):=\frac{c(D,\alpha)+\deg(\alpha)+n_{+}-2n_{-}-|D|}{2}.$
In this formula, $c(D,\alpha)$ denotes the number of circles in $D_{\alpha}$ and $|D|$ the number of components in the link represented by $D$. The function $S(D,-)$ is defined such that across a merge differential in the cube of resolutions of $D$, the increase in $\deg(\alpha)$ and the decrease in $c(D,\alpha)$ cancel out, while across a split differential, they add up to increment $S(D,\alpha)$ by $1$. As a consequence, the differentials in the refined generalized Khovanov bracket have superdegree 00. The term $n_{+}-2n_{-}-|D|$ in (32) ensures that $S(D,\alpha)$ is an integer.
We note that there would be other (inequivalent) choices for the term $n_{+}-2n_{-}-|D|$. In particular, one could replace it by ±$\pm C(D,\beta)$ for any fixed resolution $\beta$, such as the all-zero resolution. However, our choice used in (32) is more natural because it will ensure that the chain maps induced by oriented link cobordisms are homogeneous of the correct superdegree (see Remark 7 below). In this context, we point out that the quantity $\deg(\alpha)+n_{+}-2n_{-}$ in equation (32) is equal to the shift of the quantum grading mentioned at the end of the previous subsection, or to $i+w(D)$ where $i$ is the cohomological grading and $w(D)$ is the writhe.
In what follows, we will usually ignore the shifts of the supergrading, but invoke them whenever it helps simplify our proofs. Note that this approach will not limit the generality of our results because any result proven for the refined generalized Khovanov bracket can be translated back to the non-refined setting by applying a forgetful functor. In the remainder of this subsection, we will discuss some more terminology that will later be used in the refined context.
We recall that a $𝕜\Bbbk$-linear functor between two supercategories is called a superfunctor if it preserves the $\mathbb{Z}_{2}$-grading. Let $\mathcal{F}$ and $\mathcal{G}$ be two superfunctors between two supercategories $\mathcal{C}$ and $\mathcal{D}$ and $\eta=\{\eta_{x}\}_{x\in\mathcal{C}}$ be a collection of morphisms $\eta_{x}\colon\mathcal{F}(x)\rightarrow\mathcal{G}(x)$ in $\mathcal{D}$, one for each object $x\in\mathcal{C}$. Assume that all $\eta_{x}$ are homogeneous of the same $\mathbb{Z}_{2}$-degree. We will call such a collection a $\pi$-supernatural transformation if it satisfies
(33) $\eta_{y}\circ\mathcal{F}(f)=\pi^{d}\mathcal{G}(f)\circ\eta_{x}$
for every homogeneuos morphism $f\colon x\rightarrow y$ in $\mathcal{C}$ of $\mathbb{Z}_{2}$-degree $d$. Note that this terminology is slightly different from the one used in [BrundanEllis], where a supernatural transformation is assumed to split into an even and an odd component (with the two components behaving differently with respect to homogenous morphisms $f$). We note that any superfunctor $\mathcal{F}$ between two supercategories $\mathcal{C}$ and $\mathcal{D}$ extends to the refinements of these supercategories via $\mathcal{F}(x\{m\}):=\mathcal{F}(x)\{m\}$, and any $\pi$-supernatural transformation $\eta$ between two superfunctors $\mathcal{F}$ and $\mathcal{G}$ extends to a $\pi$-supernatural transformation between the extended superfunctors via $\eta_{x\{m\}}:=\pi^{m}\eta_{x}\colon\mathcal{F}(x)\{m\}\rightarrow\mathcal{G}(x)\{m\}$.
Given a supercategory $\mathcal{C}$, let $\operatorname{Kom}^{\prime}(\mathcal{C})$ denote the supercategory whose objects are chain complexes in $\mathcal{C}$ with homogeneous differentials of $\mathbb{Z}_{2}$-degree zero, and whose morphisms are (not necessarily homogenous) chain maps. Any superfunctor $\mathcal{F}$ between two supercategories $\mathcal{C}$ and $\mathcal{D}$ induces a superfunctor between the supercategories $\operatorname{Kom}^{\prime}(\mathcal{C})$ and $\operatorname{Kom}^{\prime}(\mathcal{D})$, where the induced superfunctor is given by applying $\mathcal{F}$ “termwise”. Moreover, any $\pi$-supernatural transformation $\eta$ between two superfunctors $\mathcal{F},\mathcal{G}\colon\mathcal{C}\rightarrow\mathcal{D}$ extends to a $\pi$-supernatural transformation between the induced superfunctors $\operatorname{Kom}^{\prime}(\mathcal{C})\rightarrow\operatorname{Kom}^{\prime}(\mathcal{D})$. Explicitly, let $C=(\ldots\rightarrow C^{i}\rightarrow C^{i+1}\rightarrow\ldots)$ be a chain complex in $\mathcal{C}$ with differentials of $\mathbb{Z}_{2}$-degree zero. Then the morphism $\eta_{C}$ of the extended $\pi$-supernatural transformation is given by the chain map with components $\eta_{C^{i}}\colon\mathcal{F}(C^{i})\rightarrow\mathcal{G}(C^{i})$.
@@ 3.3   Odd and Generalized Khovanov TQFTs
There is a chronological TQFT functor $F\colon\mathcal{C}hron\mathcal{C}ob^{3}_{gen}\rightarrow\mbox{Ab}$ which sends the generalized Khovanov bracket of a link diagram to the odd Khovanov complex [ORS2007, Pu2015]. This functor takes a crossingless diagram $R\subset\mathbb{R}^{2}$ to the exterior algebra
@@
where $V(R)$ is the free $\mathbb{Z}$-module formally generated by the connected components of $R$. If $S$ is a merge saddle between two components $R_{1},R_{2}\subset R$ then the map $F(S)$ is defined via the obvious quotient map $F(R)\rightarrow F(R)/(R_{1}=R_{2})$. If $S$ is a split saddle creating $R_{1},R_{2}$, then $F(S)$ is defined similarly, but via the map $F(R)/(R_{1}=R_{2})\rightarrow F(R)$ induced by left-multiplication by $R_{1}-R_{2}$ (see [ORS2007] for details).
There is a generalized TQFT functor $F_{gen}$ which takes values in the category of $𝕜\Bbbk$-modules, and which generalizes the odd TQFT functor from above and the even TQFT functor from [Kh1999]. The functor $F_{gen}$ maps the generalized Khovanov bracket to the generalized Khovanov complex (see [Pu2015]) and was originally defined in the language of [Bn2002]. We will instead define the module $F(R)_{gen}$ intrinsically by replacing the exterior algebra $F(R)=\Lambda^{*}V(R)$ by the noncommutative polynomial ring over $𝕜\Bbbk$ in the components of $R$, modulo the relations
@@
where $R_{1}$ and $R_{2}$ are components of $R$. Note that
@@
so that left-multiplication by $R_{1}+\pi R_{2}$ identifies $R_{1}$ and $R_{2}$. The map that $F_{gen}$ assigns to a split saddle can thus be defined via the map $F(R)_{gen}/(R_{1}=R_{2})\rightarrow F(R)_{gen}$ induced by left-multiplication $R_{1}+\pi R_{2}$. Here we are assuming that, near the split saddle, $R_{1}$ lies to the right of $R_{2}$ when looking in the direction of the specified orientation of the split saddle. The map that $F_{gen}$ assigns to a merge saddle is defined via the quotient map $F(R)_{gen}\rightarrow F(R)_{gen}/(R_{1}=R_{2})$, as before.
The module $F(R)_{gen}$ comes with a quantum grading and a supergrading. If $R$ has connected components $R_{1},\ldots,R_{c}$, then the quantum degree of a generator $R_{i_{1}}R_{i_{2}}\cdots R_{i_{b}}$ is given by $c-2b$, and the superdegree is given by the modulo 2 reduction of $b$. In the generalized Khovanov complex, these gradings are shifted by the shifts described in the previous subsections. One can check that the resulting shifted superdegree of a generator $g$ agrees with the modulo 2 reduction of $(j-|D|)/2$, where $j$ is the (shifted) quantum degree of $g$, and $|D|$ is the number of link components.